# From -gram to Attention: How Model Architectures Learn and Propagate Bias in Language Modeling

### v1

@@ Appendix D CrowS-Pairs Dataset
The CrowS-Pairs dataset (Nangia etΒ al., 2020) is a benchmark designed to evaluate diferent types of social biases in language models. It consists of sentence pairs, each consisted of a more stereotypical sentence (sent_more) and a less stereotypical or anti-stereotypical counterpart (sent_less). These pairs are crafted to differ minimally, focusing on variations that highlight social biases. Some sample sentences categorized by bias type is shown in TableΒ 10. The dataset encompasses nine bias types, with the distribution of examples across these categories summarized in TableΒ 9. The original dataset contains 6 fields. The description of field along with field name is as follows:
β’
@@ Appendix E Extended Result Analysis
In this section, we provide additional visualizations of our experimental results. Specifically, FiguresΒ 7, 8, and 9 display the individual bias scores across all settings. As shown in FigureΒ 7, Laplace and add- smoothing exhibit a clear staircase pattern of decreasing bias scores with increasing -gram order. In contrast, the bias scores for transformers with both soft and sparse attention do not display any consistent pattern across different settings.

### v2

@@ Appendix D CrowS-Pairs Dataset
The CrowS-Pairs dataset (Nangia et al., 2020) is a benchmark designed to evaluate different types of social biases in language models. It consists of sentence pairs, each consisted of a more stereotypical sentence (sent_more) and a less stereotypical or anti-stereotypical counterpart (sent_less). These pairs are crafted to differ minimally, focusing on variations that highlight social biases. Some sample sentences categorized by bias type is shown in Table 12. The dataset encompasses nine bias types, with the distribution of examples across these categories summarized in Table 11. The original dataset contains 6 fields. The description of field along with field name is as follows:
@@ Appendix E Extended Result Analysis
In this section, we provide additional visualizations of our experimental results. Specifically, Figures 9, 10, and 11 display the individual bias scores across all settings. As shown in Figure 9, Laplace and add- smoothing exhibit a clear staircase pattern of decreasing bias scores with increasing -gram order. In contrast, the bias scores for transformers with both soft and sparse attention do not display any consistent pattern across different settings.